OpenStudy (loser66):

Prove that every 4-cycle can be written as a product of 2-cycles Please, help

OpenStudy (loser66):

@rational @myininaya

OpenStudy (rational):

(1,2,3,4,1) = (...)(...) something like this ? if possible, can you teach me quick the definition of product in regards to groups ?

myininaya (myininaya):

I had to look this up. http://www.math.niu.edu/~jthunder/Courses/2009Fall/420/sec1/oct12/permutations2.pdf The site seems wise. I'm reading now.

OpenStudy (loser66):

ehh.... In my book, it says (1,3,2,4) can be written as (13)(12)(14)

OpenStudy (loser66):

all I have to do is prove (1324) = (13)(12)(14) but I can't get the right answer.

OpenStudy (anonymous):

@rational I found this link to be really helpful with definitions: http://people.math.sfu.ca/~jtmulhol/math302/notes/4-Permutations-CycleForm.pdf

OpenStudy (loser66):

\(\left[\begin{matrix}1&2&3&4\\3&4&2&1\end{matrix}\right]\) =\(\left[\begin{matrix}1&2&3&4\\3&2&1&4\end{matrix}\right]\)\(\left[\begin{matrix}1&2&3&4\\2&1&3&4\end{matrix}\right]\)\(\left[\begin{matrix}1&2&3&4\\4&2&3&1\end{matrix}\right]\)

OpenStudy (rational):

Exactly what I am looking for! thank you so much

OpenStudy (loser66):

That 's what I have, but I can't get them equal.

OpenStudy (loser66):

Start from far right, the last 2 terms gives me \(\left[\begin{matrix}1&2&3&4\\4&1&3&2\end{matrix}\right]\)

OpenStudy (rational):

(1,2,3,4) = (1,2,3)(1,3,4)

OpenStudy (anonymous):

\[(1,2)(1,3)(1,4)\] \[1\xrightarrow{(1,3)}3\xrightarrow{(1,2)}3\xrightarrow{(1,4)}3\\ 3\xrightarrow{(1,3)}1\xrightarrow{(1,2)}2\xrightarrow{(1,4)}2\\ 2\xrightarrow{(1,3)}2\xrightarrow{(1,2)}1\xrightarrow{(1,4)}4\\ 4\xrightarrow{(1,3)}4\xrightarrow{(1,2)}4\xrightarrow{(1,4)}1\] This gives a cycle \((1,3,2,4)\), so indeed they are equivalent.

OpenStudy (loser66):

combine with the first one from the right, I have \(\left[\begin{matrix}1&2&3&4\\4&3&1&2\end{matrix}\right]\) which is not the left hand side @rational it's the product of 3-cycle

OpenStudy (rational):

Oh the proof is about writing 4-cycle as products of 2-cycles... I see thanks :)

OpenStudy (loser66):

@SithsAndGiggles It should give the same answer when I do it under the matrix form, right? but they are not, why? It frustrates me a lot. I got confuse.

OpenStudy (loser66):

I understand the way you do. I like it too. But I have to follow my prof's way (matrix way)

OpenStudy (anonymous):

It should work with whatever notation you choose. (I'd just learned this stuff a little less than an hour ago actually, so I just went with the most economic notation :) )

OpenStudy (loser66):

Please, check my work and point out where my correct is.

OpenStudy (anonymous):

If your books says (1324)=(13)(12)(14), then that is incorrect. If we are using the standard idea of product and going from right to left, then (1324)=(14)(12)(13)

OpenStudy (loser66):

*not hihihi

OpenStudy (loser66):

@nerdguy2535 give me some link, please

OpenStudy (anonymous):

Sure, one sec.

OpenStudy (loser66):

One more thing, you see @SithsAndGiggles proof above showed that (13)(12)(14) works well. And we know that non-disjoint cycles is not commutative. --> only 1 method is a correct one. Yours or Sith??

OpenStudy (anonymous):

I was under the impression left-to-right was the standard method...

OpenStudy (anonymous):

I think we are reading this: (13)(12)(14) differently. So before we proceed lets verify something. When I see (132)(24), and someone asked me where 4 went to, I would say, "4 goes to 2, and then 2 goes to 1, so 4 goes to 1". Is that how you would read this?

OpenStudy (loser66):

yes

OpenStudy (loser66):

and it is a multiplication of 2 cycles, not a product of a non-disjoint cycle, right?

OpenStudy (anonymous):

Alright, then I stand by my statement that (1324) is not the same as (13)(12)(14), since 4 goes to 1 in the 4 cycle, and 4 goes to 2 in the product of 2 cycles. This is why your matrix multiplication isnt coming out to the right answer. You should be doing (14)(12)(13) instead.

OpenStudy (anonymous):

I'd read it as \(4\to4\), then \(4\to2\).

OpenStudy (loser66):

@nerdguy2535 I got you @SithsAndGiggles where is your question from?

OpenStudy (anonymous):

Then for you, (13)(12)(14) is correct. and you would multiply the matrices from left to right and get the right answer.

OpenStudy (anonymous):

Okay so we agree that we get two different results using two fundamentally different methods... My work suggests @Loser66 is supposed to use the left-to-right interpretation. Is that how your text/professor is teaching this?

OpenStudy (loser66):

@nerdguy2535 not just that, please, give me link, I don't want to confuse anymore. I have to master the concept.

OpenStudy (anonymous):

Here's a link to an example http://en.wikipedia.org/wiki/Symmetric_group#Multiplication

OpenStudy (loser66):

@SithsAndGiggles my prof teaches me definition, theorem, proof of theorem, and... done. I search on internet and study myself to do homework.

OpenStudy (loser66):

@nerdguy2535 I read it, it doesn't talk about decompose from left to right or right to left

OpenStudy (anonymous):

In their example they compute (13)(45)(125)(34) and get (124)(35). You could only get this if you went right to left.

OpenStudy (loser66):

I realize that some mathematicians use right to left, some use left to right. I read and confused.

OpenStudy (anonymous):

Also, I don't think we are getting two different results, I just think we are accidentally mixing up the two ways of reading these things >.<

OpenStudy (anonymous):

Like, if we had known we were taught to read these differently, there wouldnt have been a problem.

OpenStudy (loser66):

Ok, I would like to make sure again if the cycle is (12435) , then it is = (15)(13)(14)(12), right?

OpenStudy (anonymous):

Yeah. The link that myininaya posted has the general formula for decomposing a k-cycle into two cycles.

OpenStudy (anonymous):

Keep in mind that answer is not unique though. There are infinitely many ways to decompose a given cycle into two cycles.

OpenStudy (loser66):

what if it misses some, like (145) , it is = (15)(14) , right? and I can rearrange (145) as (451) , so that it is (41)(45) , right?

OpenStudy (anonymous):

Yes, that all works.

OpenStudy (loser66):

Thanks a ton. :)

OpenStudy (loser66):

@nerdguy2535 this is the link I see them decompose the product from left to right. They don't give out the procedure so that I don't know whether they proceed it from left to right or from right to left. And it confused me a lot I mean: On page 3, part Decomposition of k-cycles into 2cycles, they give example \(\alpha = (1, 3, 5)(2, 4, 7,6,8) = (1, 3)(1,5)(2, 4)(2, 7)(2, 6)(2, 8)\) Obviously, from the right hand side, they break it from left to right, right? but when I applied the rule right to left, I got wrong!!! Ha!!!